OpenStudy (anonymous):

Five more than a number tripled is half the number. What is the number? A. 6 B. –2 C. –1 D. 1

jigglypuff314 (jigglypuff314):

5 + 3x = (1/2)x solve for x...

OpenStudy (anonymous):

I believe it is B. -2 -2x3 = -6 +5 =-1 which is half of -2

jigglypuff314 (jigglypuff314):

5 + 3x = (1/2)x subtract both sides by (1/2)x -(1/2)x -(1/2)x 5 + 2.5x = 0 subtract both sides by 5 -5 -5 2.5x = -5 then divide both sides by 2.5 /2.5 /2.5 x = - 10/5 = -2 <--

OpenStudy (anonymous):

\[3x+5=\frac{ 1 }{ 2 }x\] subtract 1/2x \[\frac{ 5 }{ 2 }x+5=0\] subtract 5 \[\frac{ 5 }{ 2 }x=-5\] multiply by 2 to get rid of fraction \[5x=-10\] divide by 5 \[x=-2\]